Suppose you earn $8.50/h at a bakery. From your first paycheck you determine the $2.15/h is withheld for taxes and benefits. You work x hours during a five-day week and you spend $3.75 each day for lunch. How many hours must I work to earn $120 after taxes and benefits, and lunch expenses?
Please help. Thank you

Answer by checkley79(3341)

(8.50-2.15)x-5*3.75=120
6.35x-18.75=120
6.35x=120+18.75
6.35x=138.75
x=138.75/6.35
x=21.85 hours.
Proof:
(8.50-2.15)*21.85-5*3.75=120
6.35*21.85-18.75=120
138.75-18.75=120
120=120